OpenStudy (anonymous):

Find real numbers that satisfy the given equations |x|= 9/5 |x|= -0.8

OpenStudy (cwrw238):

|x| = 9/5 x = 9.5 , -9.5 there are no solutions to the second equation as the | | means the absolute value of x which is positive

OpenStudy (cwrw238):

* sorry the second line should be x = 9/5 , -9/5

OpenStudy (anonymous):

wow that what i as thinking.... thanks so much i really do appreciates your help.... that as two of many lol... thanks again for your time

OpenStudy (cwrw238):

yw